INAF LBC+LUCIFER queue observing, 2010 October 08-09

Observer: V. Testa, K.Boutsia, M. Fumana
Telescope Operator: D.Gonzales Huerta
Support Astronomer: O.Kuhn
Instrument: LBC - LUCIFER

Summary

good clear sky, start taking LUCIFER flats and then switch to LBC
LBC filter wheel #2 stuck (probably). had to switch instrument.
Seeing on average decent: below 1" in the IR, ~1.1-1.4 in R (according to the AG FWHM)

Data Summary

Project Objects Filter ExpTime Notes
Turatto SN2010gi J, H, Ks 40s + 80s + 80s seeing around 1"
Turatto SN2010gi 200HK 3680 s seeing below 1" - Telluric+ARC+FLAT
Longhetti S2F1-142 150Ks 4hr40min seeing below 1'' -Telluric+ARC+FLAT
PHOT_STD SA113 B,V (blue) - V,I (red) 3sec (each)  

Details

00:35 - open and prepare for IR flats (J,H,Ks)

00:55 - Ks flats (15000) 0001-0010

01:03 - H (saturates 0011-0021 saturated) flush detector back to Ks

01:05 - Ks (medium level) 0021-0030

01:07 - H (still high ~24000) 0031-0040

01:09 - back to Ks for the low ones 0041-0050
  • 4000 cts

01:11 - back to H (~15000) 0051-0060

01:13 - J (~15000) 0061-0070

01:16 - H low (~5000) 0071-0080

01:18 - J low (~4000) 0081-0090
  • Meanwhile switch on LBC before inst. switch

01:19 - SWITCH TO LBC

01:38 - Slew to ACT0219 for pointing correction
  • OB execution error - left main mirror failed to focus

01:41 Repeat ACT0219 acquisition
  • pointautoadjust/mirradjust
  • no limits hit this time...

01:47 - Slew to PHOT_STD SA113_BV_VI
  • dofpia,/First - this option introduces bigger spherical to enlarge the pupil
  • Doesn't converge - Interrupted
  • manually put Z7=-2000 (Red+Blue)
  • dofpia,/First

02:08 - Exposure STD SA113_BV_VI
  • during 4th exposure OB was blocked - lost connection with telescope
  • TCS focus error (warning)
  • LBC kill/start

02:25 - Repeat Exposure STD SA113_BV_VI

02:30 - Last Blue Image different background in lower and upper part

02:31 - Slew to AndXXI

02:35 - dofpia warning on red filters
  • redone first fastextra OB timeout error on red filter wheel #1 (r-SLOAN) hardware failure no chance to restart it.
  • turn off/turn on the instrument

02:47 - restarted, wheel #1 always blocked. Seeing not exciting. Go back to Lucifer

02:51 - SWITCH TO LUCIFER

03:25 - Slew to Hip102970 (sn2010gi) acquisition 0091-0093

03:29 - Science HIP102970
  • 5sec dit is almost saturated 0094-0095
  • redo with 2sec dit 0096-0097
  • still seems saturated - we repeat with mode NORMAL 0098-0107 (?)

03:38 - Slew to SN2010gi -acquisition- (Turatto) 108-112

03:52 - Science script for spectra SN2010gi (longslit) 113-128

05:05 - Calibration Unit IN
  • FLAT 129-159
  • ARC 158-166

05:12 - Lost azimuth drives

05:22 - Cal Unit OUT

05:23 - 2010gi acquisition for imaging 0167-0168

05:29 -Science script for imaging SN2010gi
  • H 0169-0172
  • J 0173-0176
  • Ks 0177-0180

05:41 - Slew to Telluric1 (Longhetti)
  • acquisition 0181-0184

05:59 - Science script Telluric 1 0185-0192

06:03 - Slew to Longhetti science target
  • acquisition 0193-0196

06:28 - Science script Longhetti 1st run 0197-0204
  • at the end of the OB offset Y +64 to go back to starting position

07:12 - Science script Longhetti 2nd run 0205-0212
  • at the end of the OB offset Y +64 to go back to starting position

07:57 - Science script Longhetti 3rd run 0213-0220
  • at the end of the OB offset Y +64 to go back to starting position

08:43 - Science script Longhetti 4th run 0221-0228
  • at the end of the OB offset Y +64 to go back to starting position

09:28 - Science script Longhetti 5th run 0229-0236
  • at the end of the OB offset Y +64 to go back to starting position

10:14 - Science script Longhetti 6th run 0236-0243
  • at the end of the OB offset Y +64 to go back to starting position

10:50 - End of script, guiding stopped during last image. 8th of 8 probably not good.

10:52 - Move to Telluric 2
  • acquisition 244-247

11:02 - Science script Telluric 2 0248-0255

11:07 - Back to Longhetti target
  • no star found
  • pointing correction needed (INITIALIZE)

11:39 - Science script Longhetti 7th run 0260-0267
  • at the end of the OB offset Y +64 to go back to starting position

12:24 - Science script Longhetti 8th run 0268-0270 * Stopped after the end of 3rd exposure because of twilight

12:38 - Calibration Unit IN
  • FLAT 271-560
  • ARC: Ar, Ne, Xe 561-566

Then darks and that's all folks.

-- VincenzoTesta - 09 Oct 2010
Topic revision: r7 - 11 Oct 2010, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback